OpenStudy (anonymous):

if the line contains (4, 7) and is parallel to y=5/2x + 1 how do u covert this to equation y = mx+b form

OpenStudy (smored):

first put in point-slope form (parallel= same slope) y-7=5/2(x-4) y-7=5/2*x-10 y=5/2x-3

OpenStudy (triciaal):

parallel lines have same slope using the point (4,7) 7 = 5/2(4) + b 7 = 10 + b b= -3 now you have y and b you were given m

OpenStudy (smored):

y and x has to stay as variables

OpenStudy (triciaal):

yes y =5/2 x -3

OpenStudy (smored):

how did you get -3

OpenStudy (triciaal):

look again 7 = 10 + b leave b on 1 side and keep the eqn by subtracting 10 on each side 7 - 10 = -3

OpenStudy (smored):

yes, but it is simpler by first putting it on point slope form

OpenStudy (smored):

anyways, we're both right.